OpenStudy (anonymous):

which of the following equations is parallel to the line passing through a(-2 7) b(-2 -2)

jigglypuff314 (jigglypuff314):

parallel lines have equal slope so you should first find the slope of the given points \[slope = m = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2) so your slope would be the equivalent of \[\frac{ -2 - 7 }{ -2 - (-2) }\] but you would find that you get a 0 in the denominator and you can't divide by 0 and slopes with a denominator of 0 means that it is a vertical line therefore a line parallel to a vertical line is another vertical line.